# JND-Guided Light-Weight Neural Pre-Filter for Perceptual Image Coding

### v1

@@ Abstract
Just Noticeable Distortion (JND)-guided pre-filter is a promising technique for improving the perceptual compression efficiency of image coding. However, existing methods are often computationally expensive, and the field lacks standardized benchmarks for fair comparison. To address these challenges, this paper introduces a twofold contribution. First, we develop and open-source FJNDF-Pytorch, a unified benchmark for Frequency-domain JND-Guided pre-Filters. Second, leveraging this benchmark, we propose a complete learning framework for a novel, lightweight Convolutional Neural Network (CNN). Experimental results demonstrate that our proposed method achieves better compression efficiency, consistently outperforming existing approaches across multiple datasets and encoders. In terms of computational cost, our model is exceptionally lightweight, requiring only 7.15 GFLOPs to process a 1080p image, which is merely 14.1% of the cost of a leading lightweight network in the field. Our work presents a robust solution that demonstrates superior performance and efficiency, supported by a reproducible research platform. The implementation is available at https://github.com/viplab-fudan/FJNDF-Pytorch.
@@ I Introduction
Enhancing the compression efficiency of encoders based on perceptual quality is a core challenge in perceptual coding. Among the various techniques, pre-filter guided by Just Noticeable Distortion (JND) has become a promising approach due to its portability as an encoder-only modification. Existing methods [1, 2, 3, 4, 5, 6, 7, 8, 9] can be broadly categorized by the type of JND model employed: spatial domain [1, 2, 3], frequency domain [4, 5, 6], hybrid spatial-frequency domain [7], and learning-based models [8, 9].
However, current research in this area suffers from two key limitations: 1) the majority of studies rely on Mean Opinion Score (MOS) for video quality evaluation, a metric that is time-consuming and labor-intensive, thereby hindering rapid algorithm iteration and large-scale validation; 2) many works generally fail to employ the Bjøntegaard Delta Bitrate (BD-BR) [10] to evaluate their impact on compression performance, thus lacking an objective quantification of efficiency gains.
While the few studies [5, 6] that utilize BD-BR confirm that traditional frequency-domain methods are effective, they suffer from prohibitive computational complexity. Their reliance on expensive operations like the Discrete Cosine Transform (DCT) and additional detail enhancement modules for salient regions results in a severe increase in encoding time, with overheads reaching up to 313% [5] and 123% [6].
@@
In summary, current research on JND-guided pre-filter faces two primary challenges: a limited evaluation methodology, which relies on subjective metrics and lacks standardized benchmarks for objective comparison, and the high complexity of advanced algorithms, which renders the most effective frequency-domain methods impractical for real-time application. This paper addresses these gaps with two main contributions, as illustrated in Fig. 1:
  • We develop and open-source FJNDF-Pytorch, a unified platform for the development and validation of JND pre-filter methods. The platform integrates: 1) mainstream frequency-domain JND modeling and injection methods; 2) multiple standard open-source encoders and datasets; 3) a suite of objective quality metrics, to facilitate reproducible research and address the evaluation gap.
  • We propose a complete framework for lightweight neural pre-filter. Leveraging our benchmark to generate a high-quality dataset, the framework then effectively trains a lightweight network to learn the fundamental characteristics of the reference filter via a supervised, residual learning strategy. The core of our method is a novel hybrid loss function, featuring a frequency-domain loss, which then empowers the network to surpass the performance of this training reference.
@@ III JND-Guided Lightweight Network
Our proposed method is a complete learning framework designed to train a lightweight network for JND-guided pre-filter. As illustrated in Fig. 3, we employ a supervised learning paradigm with a residual learning strategy, which reduces the task’s complexity and makes it feasible for a lightweight network. The framework is composed of three key components: a hybrid loss function designed to enable the network to surpass the performance of its reference, a data generation strategy to provide a high-quality training target, and a customized lightweight network architecture for efficient inference.
III-A Hybrid Loss Function Design
To enable the network to surpass the performance of the traditional reference algorithm, the core of our methodology is a hybrid loss function that incorporates frequency-domain priors. The total loss, ${L}_{all}$, is a weighted sum of a spatial-domain loss and our proposed frequency-domain loss, ${L}_{freq}$:
@@ III-A Hybrid Loss Function Design
The $LF$ term prevents over-smoothing by penalizing the loss of low-frequency energy, while the $HF$ term suppresses artifacts by penalizing the creation of new high-frequency energy. This constraint empowers the network to surpass the reference by correcting its frequency-domain imperfections.
III-B Data Generation Strategy
III-C Lightweight Network Architecture
We select the efficient MobileNetIE [33] as our backbone, leveraging its re-parameterization mechanism for low inference complexity. To further specialize this general-purpose model for our JND-Guided pre-filter task, we introduce two key architectural modifications. First, we replace all 5x5 convolutions with 3x3 to better align the network’s receptive field with the local nature of DCT-domain operations. Second, we design a bottleneck-expansion channel configuration, as illustrated by the Head, Body, and Tail structure in Fig. 3.

### v2

# JND-Guided Lightweight Neural Pre-Filter for Perceptual Image Coding

@@ Abstract
Just Noticeable Distortion (JND)-guided pre-filter is a promising technique for improving the perceptual compression efficiency of image coding. However, existing methods are often computationally expensive, and the field lacks standardized benchmarks for fair comparison. To address these challenges, this paper introduces a twofold contribution. First, we develop and open-source FJNDF-Pytorch, a unified benchmark for Frequency-domain JND-Guided pre-Filters. It provides a standardized environment to facilitate the rapid development and objective evaluation of pre-filter algorithms. Second, we propose a complete learning framework for neural pre-filters. Our framework trains a lightweight network to learn from a reference filter via supervised learning, with the key innovation being a frequency-domain loss that enables the network to break the performance ceiling set by its training reference. Experimental results demonstrate that our proposed method consistently outperforms existing approaches in compression efficiency across multiple datasets and encoders, while being exceptionally lightweight, requiring only 7.15 GFLOPs to process a 1080p image, which is 14.1% of a leading lightweight network. Our code is available at https://github.com/viplab-fudan/FJNDF-Pytorch.
@@ I Introduction
Enhancing the compression efficiency of encoders based on perceptual quality is a core challenge in perceptual coding. Among the various techniques, pre-filter guided by Just Noticeable Distortion (JND) has become a promising approach due to its portability as an encoder-only modification. Existing methods [1, 2, 3, 4, 5, 6, 7, 8, 9] can be broadly categorized by the type of JND model employed: spatial domain [1, 2, 3], frequency domain [4, 5, 6], hybrid spatial-frequency domain [7], and learning-based models [8, 9]. Given the superior effectiveness demonstrated by frequency-domain approaches, this paper focuses on this category of methods.
However, current research in this area suffers from three key limitations: 1) Inefficient Quality Metric: the majority of studies rely on Mean Opinion Score (MOS) for video quality evaluation, a metric that is time-consuming and labor-intensive, thereby hindering rapid algorithm iteration and large-scale validation; 2) Lack of Key Metric in Image Coding: many works generally fail to employ the Bjøntegaard Delta Bitrate (BD-BR)[10], the standard metric for quantifying efficiency gains in image coding task; and 3) High Algorithmic Complexity: While effective [5, 6], traditional frequency-domain methods suffer from prohibitive computational complexity. Their reliance on expensive operations, such modules for texture and salient region detection, introduces encoding time overheads of up to 313%[5] and 123%[6].
@@
In summary, current research on JND-guided pre-filters faces two primary challenges: limited evaluation methodology, and high complexity of advanced algorithms. This paper addresses these gaps with two main contributions, as illustrated in Fig. 1:
  • We develop and open-source FJNDF-Pytorch, a unified platform designed to facilitate the agile development and efficient validation of pre-filter methods to address the evaluation gap. The platform integrates: 1) mainstream frequency-domain JND modeling and injection methods; 2) multiple standard open-source encoders and datasets; 3) a suite of objective quality metrics.
  • We propose a complete framework for lightweight neural pre-filter. The framework trains a network to learn from a reference filter via supervised learning; however, to break the performance ceiling inherent in this approach, our key innovation is a novel frequency-domain loss. The core mechanism of this loss is the introduction of constraints that are independent of the reference target, enabling the network to correct the reference’s flaws and surpass its performance.
@@ III JND-Guided Lightweight Network
Our proposed method is a complete learning framework designed to train a lightweight network for JND-guided pre-filter. As illustrated in Fig. 3, we employ a supervised learning paradigm with a residual learning strategy, which reduces the task’s complexity and makes it feasible for a lightweight network. The framework is composed of three key components: a data generation strategy to provide a high-quality training target, a customized lightweight network architecture for efficient inference, and a hybrid loss function designed to enable the network to surpass the performance of its reference.
III-A Data Generation Strategy
III-B Lightweight Network Architecture
We select the efficient MobileIE [33] as our backbone, leveraging its re-parameterization mechanism for low inference complexity. To further specialize this general-purpose model for our JND-Guided pre-filter task, we introduce two key architectural modifications. First, we replace all 5x5 convolutions with 3x3 to better align the network’s receptive field with the local nature of DCT-domain operations. Second, we design a bottleneck-expansion channel configuration, as illustrated by the Head, Body, and Tail structure in Fig. 3.
III-C Hybrid Loss Function Design
To enable the network to surpass the performance of the traditional reference algorithm, the core of our methodology is a hybrid loss function that incorporates spatial and frequency loss. The total loss, ${L}_{all}$, is a weighted sum of a spatial-domain loss and our proposed frequency-domain loss, ${L}_{freq}$:
@@ III-C Hybrid Loss Function Design
Specifically, the $LF$ term penalizes the network when the magnitude of its low-frequency coefficients is less than the original’s, while the $HF$ term applies a penalty when the magnitude of its high-frequency coefficients exceeds the original’s.